OpenStudy (help108):

MEDAL!!!!!!! Which statement is true about the equation 4.5z = 4.5z − 4.25 + 6.25? It has no solution. It has one solution. It has two solutions. It has infinitely many solutions.

OpenStudy (help108):

@AloneS help plzzz

OpenStudy (help108):

hello? @AloneS

alones (alones):

Okay 4.5z=4.5z−4.25+6.25 simplify both sides of the equation. 4.5z=4.5z−4.25+6.25 4.5z=4.5z+2 subtract 4.5z from both sides. 4.5z−4.5z=4.5z+2−4.5z 0=2 So far i got no solution

OpenStudy (help108):

so is this your final anwser no solution or theirs still more to go?

alones (alones):

No solution is final answer

OpenStudy (help108):

thathanks